Solve for the net intensity at point A.

   q1              9 m                                      A         2 m                   q2

I

 . _______________________________O__________________________.

-5 µC                                                                                             + 3 µC


1
Expert's answer
2021-02-16T10:30:18-0500

Answer

Net electric field intensity at point a can be given as

"E=\frac{kq_1}{r_1^2}+\frac{kq_2}{r_2^2}\\E=k(\frac{q_1}{r_1^2}+\frac{q_2}{r_2^2})"


By putting value


"E=(9\times10^9)(\frac{5\times10^{-6}}{9^2}+\frac{3\times10^{-6}}{2^2})"


"E=7.3\times10^{-3}N/c"